\vspace{6mm}

\begin{thebibliography}{36}


\bibitem{Bateman}
{\bf H.Bateman, A.Erdelyi, W.Magnus, F.Oberhettinger, F.G.Tricomi.} {\it Higer transcendental functions, }
Volume II,  California Institute of Technology. Mc.Graw - Hill Book Company INC., USA, 1953;
Renewed 1981.

\bibitem{Bertin1}
{\bf Bertin K.}  (2004). {\it Asymptotically exact minimax estimation in sup-norm for anisotropic
H\"older classes.} Bernoulli,  {\bf 10,}  873 \ - \ 888. MR2093615

\bibitem{Bertin2}
{\bf Bertin K.} (2004). {\it Estimation asymptotiquement exacte en norme sup de fonctions multidimensionnelles.}
Ph.D. thesis, Universiti  Paris 6.


\bibitem{Charpentier}
{\bf A Charpentier and E Flachaire. } {\it Log-transform kernel density estimation of income distribution.}
L'Actualite Economique, {\bf 91;} 141 \ - \ 159, 2015.

\bibitem{Compte}
{\bf Fabienne Compte and Nicolas Marie.} {\it  Bandwidth selections for the Wolverton - Wagner estimator.}
arXiv:1902.00734v2 [math.ST] 12 Oct 2019


\bibitem{Courant Hilbert}
{\bf Courant, R; Hilbert, D.}  (1953). Methods of Mathematical Physics. I (First English ed.). New York: Interscience
Publishers, Inc.


\bibitem{Devroe}
{\bf Devroye L.}  (1979)  {\it On the pointwise and integral convergence of recursive kernel estimates of probability
densities.}  Util. Math., {\bf  15,} 113 \ - \ 128.

\bibitem{Devroe Giorfi}
{\bf Devroye Giorfi.}   {\it Nonparametric density estimation; the L1 view.}
New York : John Wiley, 1985.

\bibitem{Epanechnikov}
{\bf V. A. Epanechnikov.} {\it Nonparametric estimation
of a multidimensional probability density.} Teor. Veroyatnost.
i Primenen., 1969, Volume 14, Issue 1, 156 \ - \ 161, (in Russian).

\bibitem{Gasser}
{\bf Gasser, T. and M\"uller, H. G.}  (1984).  {\it Estimating regression functions and their derivatives by the kernel method.}
 Scandinavian Journal of Statistics, {\bf 11},  171 \ - \ 185.

\bibitem{Gelfand Fomin}
{\bf Gelfand, I. M.; Fomin, S. V.}  {\it Calculus of variations.}   Mineola, New York: Dover Publications. p. 3. ISBN 978-0486414485.
 (2000). Silverman, Richard A. (ed.).


\bibitem{Goldenshluger Lepski}
{\bf A.Goldenshluger, and O.Lepski.}   {\it Bandwidth selection in kernel density estimation: oracle
inequalities and adaptive minimax optimality.} The Annals of Statistics, {\bf 39,}  1608 \ - \ 1632, 2011.


\bibitem{Hansen}
{\bf Bruce E. Hansen.} {\it Lecture Notes on Nonparametrics.}
University of Wisconsin, Spring, 2009


\bibitem{Hardle}
{\bf H\"ardle Wolfgang.} {\it Applied Nonparametric Regression.}
Humboldt-Universitat zu Berlin Wirtschafts wissenschaftliche Fakultat Institut f\"ur Statistik und Okonometrie; Spandauer Str.
 1D–10178, Berlin, 1994.


\bibitem{Ibragimov Khasminskii 1}
{\bf Ibragimov, I. A. and Khasminskii, R. Z.} (1980). {\it An estimate of the density of a distribution. } Zap. Nauchn. Sem.
Leningrad. Otdel. Mat. Inst. Steklov. (LOMI), {\bf 98,}  61 \ - \ 85, (in Russian).

\bibitem{Ibragimov Khasminskii 2}
{\bf Ibragimov, I. A.  and Khasminskii, R. Z.} (1981). {\it More on estimation of the density of a distribution.} Zap. Nauchn.Sem. Leningrad.
Otdel. Mat. Inst. Steklov. (LOMI), {\bf 108,} 72 \ - \ 88,  (in Russian).


\bibitem{Khardani}
{\bf Salah Khardani, Yousri Slaoui.}  {\it Recursive Kernel Density Estimation and Optimal Bandwidth Selection Under $ \ \alpha \ $  Mixing Data.}
Journal of Statistical Theory and Practice,  (2019),  {\bf 13:36,} https://doi.org/10.1007/s42519-018-0031-61, JOURIGINAL ARTICLE.

\bibitem{Lerasle}
{\bf M. Lerasle, N. Magalhaes and P. Reynaud-Bouret.}  {\it Optimal Kernel Selection for Density
Estimation.}  High-Dimensional Probability VII: The Cargese Volume, Prog. Probab. {\bf 71,}
Birkhauser, 425 \ - \ 460, 2016.

\bibitem{Nadaraya}
{\bf Elizbar Nadaraya, Petre Babilua.} {\it On the Wolverton \ - \ Wagner Estimate of a Distribution Density.}
BULLETIN OF THE GEORGIAN NATIONAL ACADEMY OF SCIENCES, 175, {\bf  ½1,} 2007,  Mathematics.


\bibitem{Parzen}
{\bf Parzen E.}  (1962.) {\it On estimation of a probability density and mode.}  Ann. Math. Stat., {\bf  33. } 1065 \ - \ 1076.

\bibitem{Rosenblatt}
{\bf Rosenblatt  M.}   (1956.)  {\it Remarks  on  some  nonparametric  estimates  of  a  density  function.}   Ann. Math. Stat., {\bf 27,} 832 \ - \ 837.

\bibitem{Simone}
{\bf Andrea De Simone, Alessandro Morandinia.}   {\it Nonparametric Density Estimation from Markov Chains.} \\
arXiv:2009.03937v1 [stat.ME] 8 Sep 2020


\bibitem{Slaoui}
{\bf Slaoui  Y.}   (2013.)  {\it Large  and  moderate  principles  for  recursive  kernel  density  estimators  defined  by  stochastic
approximation method.} Serdica Math. J., {\bf 39}, 53 \ - \ 82.

\bibitem{Stute 1}
{\bf Stute, W.} (1982). {\it A law of the logarithm for kernel density estimators.}  Ann. Probab., {\bf 10,}
414 \ - \ 422. MR647513

\bibitem{Stute 2}
{\bf Stute, W.}  (1984). {\it The oscillation behavior of empirical processes: The multivariate case.}
Ann. Probab., {\bf 12,} 361 \ - \ 379. MR735843

\bibitem{Tsybakov1}
{\bf Tsybakov A.B.}  (1990).  {\it Recurrent estimation of the mode of a multidimensional distribution.}  Probl. of  Inf. Transm.,
{\bf 8,} 119 \ - \ 126.


\bibitem{Tsybakov2}
{\bf A.B.Tsybakov.}  {\it Introduction to Nonparametric Estimation.} Springer, 2009.

\bibitem{Wegman}
{\bf E.J.Wegman and H.I.Davies.}  {\it Remarks on Some Recursive Estimators of a Probability
Density.}  The Annals of Statistics, {\bf 7,}   316 \ - \ 327, \ 1979.

\bibitem{Wolverton Wagner}
{\bf Wolverton C, Wagner T.J.}  (1969). {\it Asymptotically optimal discriminant functions for pattern classification.}
 IEEE Trans Inform Theory,  {\bf 15,} 258 \ - \ 265.

\end{thebibliography}

\documentstyle[12pt]{article}

  \newcommand{\const}{\rm const}
  \newcommand{\Var}{\rm Var}

\begin{document}

 \begin{center}

 \ {\bf Signed variable  optimal kernel for} \par

 \vspace{4mm}

 {\bf non - parametric density estimation. }\par

\vspace{5mm}

{\bf  M.R.Formica, E.Ostrovsky, and L.Sirota. } \par

 \vspace{5mm}

 \end{center}

 \ Universit\`{a} degli Studi di Napoli Parthenope, via Generale Parisi 13, Palazzo Pacanowsky, 80132,
Napoli, Italy. \\

e-mail: mara.formica@uniparthenope.it \\

\vspace{3mm}

Department of Mathematics and Statistics, Bar-Ilan University,
59200, Ramat Gan, Israel. \\

\vspace{4mm}

e-mail: eugostrovsky@list.ru\\
Department of Mathematics and Statistics, Bar-Ilan University,\\
59200, Ramat Gan, Israel.

\vspace{4mm}

e-mail: sirota3@bezeqint.net \\

\vspace{5mm}

\begin{center}

 \ {\bf Abstract.}

 \end{center}

  \ We derive the optimal signed variable in general case kernels for the classical statistic density estimation, which are some
 generalization of the famous  Epanechnikov's  ones.\par

 \vspace{5mm}

 \ {\it Key words and phrases.} Probability,  random variable and vector (r.v.),  density of distribution, H\"older's and other
 functional class of functions,  kernel, optimization, Lagrange's factor method, Legendre's ordinary and dilated polynomials,
  Euler's equation,  even function, fractional order, examples,
Epanechnikov's kernel and  generalized Epanechnikov's kernel (GEK),  bandwidth, conditions of orthogonality, bias, variance, expectation,
Parzen - Rosenblatt and  recursive  Wolverton - Wagner  statistical density estimation. \par

\vspace{5mm}

\section{Statement of problem. Notations and definitions. Previous results.}

\vspace{5mm}

 \ Let  $ \ (\Omega,  M, {\bf P})  \ $  be probability space with expectation $ \ {\bf E} \ $ and variance $ \ \Var. \ $
 \ Let also  $ \ \{\xi_k \}, \ k = 1,2,\ldots,n \ $ be a sequence of independent, identical distributed (i, i.d.) random variables (r.v.)
 taking the values in the  real axis $ \ R, \ $ and having certain {\it non - known} density of a distribution $ \ f = f(x), \ x \in R. \ $ \par
  \ We suppose further that this function belongs to the certain space of all the numerical valued $ \  2m \ -  \ $ times continuous bounded differentiable
 functions  $ \ C^{2m}(R), \ m = 1,2,\ldots,   \ $ having a finite norm

\begin{equation}  \label{C2m space}
||f||_{2m} \stackrel{def}{=} \max_{k= 0,1,2,\ldots,2m} \ \sup_{x \in R} | f^{(k)}(x) | < \infty.
\end{equation}

\  Let also $ \ K = K(x), \ x \in  \ $ be certain {\it kernel},  i.e. measurable {\it even} function having finite support:

\begin{equation} \label{theta support}
\exists \theta \in (0,\infty) \ \forall x: |x| > \theta \Rightarrow K(x) = 0,
\end{equation}

 for which

\begin{equation} \label{Kernel}
\int_{R} K(x) \ dx = 1.
\end{equation}

 \ We impose also the following conditions on this kernel.

\begin{equation} \label{K cond 1}
K(-x) = K(x);  \ V_2(K) := \int_{R} K^2(x) dx < \infty;
\end{equation}

\begin{equation} \label{K cond 2}
K(\cdot) \in C(R), \ \int_{R} |K(x)| dx < \infty.
\end{equation}

 \ The following conditions, which are also to be presumed, may be named as  {\it conditions of orthogonality:}

\begin{equation} \label{first cond orthog}
\forall l = 1,2,\ldots,2m - 1 \ \Rightarrow \int_{R} x^l \ K(x) \ dx = 0.
\end{equation}

 \ Recall that the classical Parzen - Rosenblatt estimation $ \ f_n(x) = f_n^{PR}(x)  \ $
 of the density function $ \ f(x) \ $ has a form

\begin{equation} \label{OPR}
f_n(x) \stackrel{def}{=} \frac{1}{nh} \ \sum_{i=1}^n K \left( \ \frac{x - \xi_i}{h}  \ \right).
\end{equation}
 \ Here $ \ h = h(n)  \ $  be a deterministic positive sequence such that $ \ \lim_{n \to \infty} h(n) = 0 \ $
 and $\ \lim_{n \to \infty} n h(n) = \infty, \ $  see \cite{Parzen}, \cite{Rosenblatt}. \par

 \ These and alike estimations  was study in many works, see e.g. \cite{Devroe},  \cite{Epanechnikov},  \cite{Ibragimov Khasminskii 1},
 \cite{Ibragimov Khasminskii 2},  \cite{Nadaraya}, \cite{Parzen}, \cite{Rosenblatt},
 \cite{Slaoui},  \cite{Tsybakov1}, \cite{Tsybakov2}, \cite{Wegman}, \cite{Wolverton Wagner} etc. The optimal choose of
$ \ h = h(n)  \ $ and the kernel $ \ K(x) \ $ are devoted the following works
\cite{Compte},   \cite{Goldenshluger Lepski},  \cite{Khardani}, \cite{Lerasle}. The case when the r.v. - s. $ \ \{ \ \xi_i \ \} \ $
 are (weakly) dependent is investigated in  \cite{Khardani}, \cite{Simone}. \par

\  The conditions (\ref{first cond orthog})  may be used only for the investigation of {\it bias} $ \ \delta_n(x) \ $ of these statistics.
Namely, denote

$$
\delta_n(x) \stackrel{def}{=} {\bf E} f_n(x) - f(x);
$$
then under our conditions

\begin{equation} \label{bias est}
|\delta_n(x)| \le C_1(f) h^{2m}(n).
\end{equation}

 \ The variation of $ \ f_n(x) \ $ may be estimated as follows

\begin{equation} \label{Var es}
\Var \{ \ f_n(x) \ \} \le C_2(f) \ n^{-1} \ \int_{-\infty}^{\infty} K^2 (y) \ dy.
\end{equation}

\vspace{4mm}

\begin{center}

{\sc Statement of an optimization problem.}

\end{center}

\vspace{4mm}

 \ The relations   (\ref{Var es}) common with the limitations  (\ref{theta support}),   (\ref{Kernel}), (\ref{K cond 1}), (\ref{K cond 2}),
(\ref{first cond orthog}) lead as was shown  by V.A.Epanechnikov  in \cite{Epanechnikov} to the following setting of the constrained  optimization problem:

\begin{equation} \label{statement}
\int_{-\infty}^{\infty} K^2(y) \ dy \to \min,
\end{equation}

\vspace{3mm}

under limitations

\vspace{3mm}

\begin{equation} \label{restr1}
\int_{-\infty}^{\infty} K(y) \ dy = 1; \ \forall l=1,2,3, \ldots,2m - 1 \ \Rightarrow \int_{-\infty}^{\infty} y^l \ K(y) \ dy = 0;
\end{equation}

\begin{equation} \label{2m moment}
\int_{-\infty}^{\infty} y^{2m} K(y) dy = 1,
\end{equation}

\begin{equation} \label{theta restr}
\exists \theta \in (0,\infty) \  \forall y: |y| > \theta \  \Rightarrow K(y) = 0.
\end{equation}

 \vspace{5mm}

 \ This problem was solved in the case $ \ m = 1 \ $ by V.A.Epanechnikov in \cite{Epanechnikov}, 1969 year: $ \ \theta = \sqrt{5} \ $ and

\begin{equation} \label{Epanechnikov solution}
|y| \le \sqrt{5} \Rightarrow K(y) = \frac{3}{4 \sqrt{5}} - \frac{3y^2}{20 \sqrt{5}}.
\end{equation}

\vspace{4mm}

 \  {\bf  Our aim in this short report is to find the optimal kernel $ \ K \ $ for arbitrary natural value } $ \ m = 2,3,\ldots.\ $ \par
{\bf The case when the number $ \ m \ $ is fractional, will be considered the fourth section.} \par

\vspace{5mm}

 \ Another motivation of these statement of problem appears from the famous result belonging to W.Stute \cite{Stute 1}, \cite{Stute 2}:

$$
\lim_{n \to \infty} \sqrt{ \ \frac{n h(n)}{2 |\ln h(n)|} \ }  ||f_n - {\bf E} f_n||_{\infty} = ||f||_{\infty}^{1/2} \cdot V_2(K),
$$
where as ordinary $ \ ||g||_{\infty} = \sup_{x \in R} |g(x)|. \ $\par

\vspace{4mm}

 \ {\bf Remark 1.1.} When $ \ m \ge 2, \ $ we imposed in particular the following condition on the kernel $ \ K(\cdot): \ $

$$
\int_{-\theta}^{\theta} y^2 \ K(y) \ dy = 0.
$$
 \ Therefore, the kernel $ \ K(\cdot) \ $ can not take only non - negative values. \par

\vspace{5mm}

\section{Main result.}

\vspace{5mm}

\begin{center}

 {\sc Some facts  about Legendre's polynomials.    }

\end{center}

  \ The classical Legendre's  polynomials with support on the closed interval $ \ X := [-1,1], \ $ denotes as ordinary by
$ \ P_k(x), \ x \in [-1,1], \ k = 0,1,2,\ldots  \ $ may be defined for instance as follows

$$
P_k(x) = \frac{1}{2^k \ k!} \ \frac{d^k}{d x^k} (x^2 - 1)^k,
$$
Rodriguez'formula. The polynomial $ \ P_k(x) \ $ is really polynomial of  degree $ \ k. \ $   These polynomials are orthogonal:

$$
\int_{-1}^1 P_k(x) \ P_l(x) \ dx  = \frac{2}{2k +1} \ \delta_{k,l}, \ k,l = 0,1,2,\ldots,
$$
where $ \ \delta_{k,l} \ $ is Kroneker's symbol. Following,

\begin{equation} \label{orthogon}
\forall k \ge 1, l < k, \ l \ge 0 \ \Rightarrow \int_{-1}^1 P_k(x) \ x^l \ dx = 0.
\end{equation}
\ Many properties of these polynomials may be found, e.g.  in  the classical book \cite{Bateman}. We will use
the following relations:

\begin{equation} \label{value at one}
\forall k = 0,1,2,\ldots \ P_k(1) = 1;
\end{equation}

\begin{equation} \label{import int}
 2\mu(k) \stackrel{def}{=} \int_{-1}^1 x^k \ P_k(x) \ dx = \frac{2^{k+1} \ (k!)^2}{(2k + 1)!}.
\end{equation}

 \ Some examples: $ \ P_0(x) = 1,\ P_1(x) = x, \ P_2(x) = 0.5 (3x^2 - 1), \ $

\begin{equation} \label{P4}
 P_4(x) = 8^{-1} (35  x^4 - 30 x^2 +3).
\end{equation}

\vspace{4mm}

 \ {\bf Definition 2.1.} Let $ \ \theta = \const \in (0,\infty). \ $ The {\it dilated} Legendre's polynomial
 $ \ L_k^{\theta}(x), \ k=0,1,2,\ldots; \ x \in [-\theta,\theta] \ $ of degree $ \ k, \ k = 0,1,2,\ldots \ $
 with parameter $ \ \theta \ $  is defined as follows.

\begin{equation} \label{dilated pol}
L_k^{\theta}(x) \stackrel{def}{=} \frac{1}{\theta} \ P_k \left( \ \frac{x}{\theta}  \ \right).
\end{equation}

\vspace{3mm}

 \ Of course, the properties of these polynomials  follows from ones for Legendre's polynomials. For instance,
$ \ L_{2s}(\cdot), s = 0,1,2,\ldots \ $ is even function,  $ \ L_{2s+1}(\cdot), s = 0,1,2,\ldots \ $ is odd;  \\
relations of orthogonality

\begin{equation} \label{ortho dilat}
\int_{-\theta}^{\theta} x^l \ L_k^{\theta}(x) \ dx = 0, \ k > l, \ l = 0,1,\ldots, k-1;
\end{equation}
 as well as

$$
 L_k^{0}(0) = 1/\theta = L_k(\theta),
$$

\begin{equation} \label{L import int}
\int_{-\theta}^{\theta} y^k \ L_k^{\theta}(y) \ dy =  \theta^k \cdot \frac{2^{k+1} \ (k!)^2}{(2k + 1)!},
\end{equation}

\begin{equation} \label{L orthogon}
\int_{-\theta}^{\theta} L_k(y) \ L_l(y) \ dy = \frac{2/\theta}{2k + 1} \ \delta_{k,l}.
\end{equation}
\ In particular,

\begin{equation} \label{L norming}
\int_{-\theta}^{\theta} L_k^2(y) \ dy = \frac{2/\theta}{2k + 1}.
\end{equation}

 \vspace{4mm}

  \ Let us return to the formulated above optimization problem  (\ref{statement}), (\ref{restr1}), (\ref{2m moment}), (\ref{theta restr}). \par

\vspace{4mm}

 \ {\bf Theorem 2.1.} The formulated optimization problem has an unique solution $ \ K_0(y) \ $ and having a
 following form

\begin{equation} \label{optimal 2m kernel}
K_0(y) = \frac{1}{2 \theta_0} - \frac{1}{2} L^{\theta_0}_{2m}(y) =  \frac{1}{2 \theta_0} -  \frac{1}{2 \theta_0} P_{2m} \left( \ \frac{y}{\theta_0}  \ \right),
\end{equation}

when $ \ |y| \le \theta_0, \ $ and $ \ K_0(y) = 0 \ $ otherwise. Here

\begin{equation} \label{theta 0}
\theta_0 = \left[ \  \frac{1}{1 -  \mu(2m)} \ \right]^{1/(2m)},
\end{equation}

\vspace{3mm}

and wherein under our restrictions on the kernel $ \ K(\cdot) \ $

\begin{equation} \label{minimal value}
\min \int_R K^2(y) dy =  \int_R K^2_0(y) \ dy  = \frac{1}{2 \theta_0} \cdot \frac{4m+3}{4m+1}.
\end{equation}

\vspace{3mm}

 \ Recall that

$$
\mu(2m) = \frac{2^{2m} \ (2m)!}{(4 m + 1)!}.
$$

\vspace{4mm}

 \ {\bf Proof.} We will follow the V.A.Epanechnikov,  author of the article \cite{Epanechnikov}, applied the famous
 Legendre factors method. The Euler's equation for this problem give us the following equality for the optimal kernel $ \ K(\cdot) \ $

 $$
 K(y) = \sum_{s=0}^{2m} \lambda_s \ y^s, \ |y| \le \theta.
 $$
 \ On the other words, $ \ K \ $ is polynomial of degree $ \ \le 2m, \ $  inside the interval $ \ [-\theta,\theta]. \ $
As long as the kernel is even function,

$$
\lambda_{2 r + 1} = 0, \ r = 0,1,\ldots, m - 1.
$$
 \ Further, it follows from the conditions of orthogonality that also

$$
\lambda_2 = \lambda_4 = \ldots \lambda_{2m - 2} = 0.
$$
 \ Therefore, the optimal kernel has a form

$$
K(y) = a - b L^{\theta}_{2m}(y) = a - \frac{b}{\theta} P_{2m} \left( \ \frac{y}{\theta} \ \right), \ a,b = \const.
$$
 \ We deduce substituting  the value $ \ y = \theta \ $ and taking into account the relations $ \ K(\theta) = 0 \ $
 and $ \ P_{2m}(1) = 1  \ $

\begin{equation} \label{first eq}
a = \frac{b}{\theta}.
\end{equation}

 \ Secondly,

 $$
 1 = \int_{-\theta}^{\theta} K(y) dy = 2 a \theta,
 $$
following $ \ a = 1/(2\theta) \ $ and hence $ \ b = 1/2, \ $ so that

$$
K(y) = \frac{1}{2 \theta} - \frac{1}{2} L_{2m}^{\theta} (y) = \frac{1}{2 \theta} - \frac{1}{2 \theta} P_{2m} \left( \ \frac{y}{\theta}  \ \right).
$$

\vspace{3mm}

 \ Thirdly,

$$
 1 = \int_{-\theta}^{\theta} y^{2m} \ K(y) \ dy  = \theta^{2m} - 0.5 \ \theta^{2m} \int_{-1}^1 z^{2m} \ P_{2m}(z) \ dz =
$$

$$
\theta^{2m} - \theta^{2m} \ \mu(2m).
$$

\vspace{3mm}

 \ Now, proposition (\ref{theta 0}) there holds, as well. Thus,

$$
K_0(y)= \frac{1}{2 \theta_0} -  \frac{1}{2 \theta_0} P_{2m} \left( \ \frac{y}{\theta_0}  \ \right), \ |y| \le \theta_0,
$$

 \ Ultimately, the equality (\ref{minimal value}) follows immediately from ones (\ref{L orthogon}) and (\ref{L norming}).\par

\vspace{3mm}

 \ This completes the proof of our theorem. \par

\vspace{5mm}

 \ {\bf Examples.} Of course, for the case $ \ m = 1 \ $ we obtain the classical Epanechnikov's kernel. \par
  \ Let now $ \ m = 2. \ $ We deduce after some calculations

$$
\theta_0 = \left[ \ {\frac{63}{11}} \ \right]^{1/4},
$$
and correspondingly

$$
K_0(y) = \frac{1}{2\theta_0} \ \left\{ \ 1 - \frac{1}{8} \left( \ 35 y^4 \theta_0^{-4} -30 y^2 \theta_0^{-2} +3 \ \right)  \ \right\}, \
|y| \le \theta_0,
$$
 and of course

$$
K_0(y) =0, \ |y| > \theta_0.
$$

 \ Wherein

\begin{equation} \label{minimal part  value}
\min \int_R K^2(y) \ dy =  \int_R K^2_0(y) \ dy  = \frac{11}{18 \ \theta_0}.
\end{equation}

 \vspace{5mm}

\section{The case of fractional order.}

 \vspace{5mm}

 \ Let $ \ \beta \ $ be arbitrary  {\it  fractional positive} number; denote by $ \ l = [\beta] \ $ its positive integer part.
We suppose that the function $ \ f(\cdot) \ $ belongs to the space $ \ \Sigma(\beta). \ $ This imply that all the derivatives
$ \ f^{(k)}(x), k = 0,1,\ldots,l \ $ are continuous and bounded and that the last continuous derivative $ \ f^{(l)}(x) \ $ is
bounded  and satisfies the H\"older's condition with power $ \ \beta - l. \ $ \par
 \ We impose on the kernel $ \ K(\cdot) \ $ conditions alike ones in the first section:  $ \ K = K(x), \ x \in  \ $ be certain
 even function having finite support:

\begin{equation} \label{theta beta support}
\exists \theta \in (0,\infty) \ \forall x: |x| > \theta \Rightarrow K(x) = 0,
\end{equation}

 for which

\begin{equation} \label{Kernel beta}
\int_{R} K(x) \ dx = 1,
\end{equation}

\begin{equation} \label{K cond  beta 1}
K(-x) = K(x); \ \int_{R} K^2(x) dx < \infty;
\end{equation}

\begin{equation} \label{K  cond beta 2}
K(\cdot) \in C(R), \ \int_{R} |K(x)| dx < \infty.
\end{equation}

 \ The following conditions may be named as  before {\it conditions of orthogonality:}

\begin{equation} \label{cond orthog}
\forall r = 1,2,\ldots, l \ \Rightarrow \int_{R} x^r \ K(x) \ dx = 0.
\end{equation}

 \ Ultimately,  suppose

\begin{equation} \label{beta mom}
\int_R |x|^{\beta} K(x) \ dx < \infty.
\end{equation}

 \ It is known that under these conditions the bias $ \ \delta_n(x) \ $ of Parzen - Rosenblatt's estimation  $ \ f_n(x), \ $  as well as
 for Wolverton - Wagner's  ones  obey's a following  property

$$
\sup_x |\delta_n(x)| = \sup_x | \ {\bf E} f_n(x) - f(x) \ | \le C(\beta,f) h_n^{\beta},
$$
see e.g. \cite{Compte},  \cite{Ibragimov Khasminskii 1}, \cite{Ibragimov Khasminskii 2}. \par
 \ We get following again V.A.Epanechnikov \cite{Epanechnikov} the following variational statement of problem
under formulated above in this section restrictions

\begin{equation} \label{beta optimiz}
\int_R K^2(y) \ dy \to \min_K
\end{equation}
where {\it in addition} $ \ \int_R |y|^{\beta} \ K(y) \ dy = 1.\ $ \par

\vspace{4mm}

 \ {\bf Theorem 3.1.} The (unique)  solution of this problem has a form

\begin{equation} \label{theta0 beta}
 \theta_0 = (2 \beta + 1)^{1/\beta};
\end{equation}

\begin{equation} \label{K0 beta}
K_0(y) = \lambda - \mu |y|^{\beta}, \ |y| \le \theta_0, \ K_0(y) = 0, \ |y| > \theta_0;
\end{equation}

\begin{equation} \label{lambda beta}
\lambda =  \frac{\beta + 1}{2\beta} \cdot (2 \beta + 1)^{-1/\beta},
\end{equation}

\begin{equation} \label{lambda beta}
\mu = \frac{\beta + 1}{2\beta} \cdot (2 \beta + 1)^{-(\beta + 1)/\beta},
\end{equation}

and wherein under our condition

\begin{equation} \label{minimal value 1}
\min \int_R K^2(y) \ dy = \int_R K_0^2(y) \ dy = \mu^2 \ \theta_0^{2 \beta + 1} \ \cdot \left\{ \ \frac{2 \beta^2}{(2 \beta + 1)(\beta +1)}  \ \right\}=
\end{equation}

\begin{equation} \label{minimal value 2}
(\beta +1) \cdot (2 \beta + 1)^{- (\beta + 1)/\beta}.
\end{equation}

\vspace{4mm}

 \ {\bf Proof.} The Euler's equations  for this problem give us as before the following form for the
optimal kernel

\begin{equation} \label{form of K}
K(y) = \lambda - \mu |y|^{\beta} + \sum_{j=1}^l  \nu_j y^j, \ |y| \le \theta.
\end{equation}
 \ Since the function $ \ K(\cdot) \ $ is even, $ \ \nu_1 = \nu_3 = \ldots  = 0.  \ $  Further, it follows from the
relations of orthogonality that all the other coefficients  $ \ \nu_j \ $ are absent; so the optimal kernel $ \ K \ $
has a form inside the closed interval $ \ y \in [ - \theta,\theta]: \  $

\begin{equation} \label{form of K}
K(y) = \lambda - \mu \ |y|^{\beta}.
\end{equation}

\ As long as $ \ K(\theta) = 0, \ $

\begin{equation} \label{first eq}
\lambda = \mu \cdot \theta^{\beta}.
\end{equation}

\vspace{3mm}

\ Secondly,

$$
1 = \int_{-\theta}^{\theta} K(y) \ dy =  2 \left[ \ \lambda \theta - \mu \frac{\theta^{\beta}}{\beta + 1}   \ \right],
$$

following

\begin{equation} \label{second eq}
\lambda \theta -   \mu \cdot \frac{\theta^{\beta}}{\beta + 1}  = \frac{1}{2}.
\end{equation}

\vspace{3mm}

 \ Thirdly,

$$
1 = \int_{-\theta}^{\theta} |y|^{\beta} K(y) \ dy =  2 \ \lambda \frac{\theta^{\beta}}{\beta + 1} - 2 \ \mu \frac{\theta^{2 \beta + 1}}{2 \beta + 1},
$$
 or equally

\begin{equation} \label{third eq}
\lambda \frac{\theta^{\beta + 1}}{\beta + 1} - \mu \frac{\theta^{2 \beta +1}}{2 \beta + 1}  = \frac{1}{2}.
\end{equation}

\vspace{3mm}

 \ Solving  the system of equations (\ref{first eq}), \  (\ref{second eq}) and  (\ref{third eq}), we obtain the assertion of theorem 3.1. \par

\vspace{3mm}

 \ The equalities  (\ref{minimal value 1}), (\ref{minimal value 2}) may be obtained after simple calculations.\par

\vspace{3mm}

 \ When for instance $ \ \beta = 3/2, \ $ then

$$
\theta_0 = 4^{2/3}, \ \lambda = \frac{5}{6} \cdot 4^{-2/3}, \ \mu =   \frac{5}{6} \cdot 4^{-5/3},
$$

and

$$
\min \int_R K_0^2(y) \ dy = 5 \cdot 2^{-1/3}.
$$

\vspace{5mm}

 \ {\bf Remark 3.0.} \par

  \ Note that in the case of fractional value $ \ \beta \ $ the optimal kernel $ \ K_0 \ $ is non - negative!

\vspace{5mm}

 \ {\bf Remark 3.1.} \par

 \ It is  interest to note that if we  choose as the value $ \ \beta \ $ an {\it integer} value $ \ \beta := 2, \ $ we obtain the classical
 Epanechnikov's kernel

$$
\theta_0 = \sqrt{5}, \  \lambda = \frac{3}{4 \sqrt{5}}, \ \mu = \frac{3}{20 \sqrt{5}}.
$$

 \ In this case the minimal value of $ \ \int_R K_0^2(y) \ dy \ $ is equal to $ \ 3/(5 \sqrt{5}). \ $

\vspace{4mm}

\ {\bf Remark 3.2.} We do not use in this section, i.e. in the case of fractional value of the parameter $ \ \beta, \ $
the theory of Legendre's polynomials, in contradiction to the foregoing sections. \par

\vspace{5mm}

\section{Another statement of problem.}

 \vspace{5mm}

 \ We retain all the restrictions and conditions of the foregoing section.  Denote  in addition

$$
J_{\beta} = J_{\beta}(K) := \int_R |y|^{\beta} \ K(y) \ dy; \hspace{4mm}  V_2 = V_2(K) := \int_R K^2(y) \ dy,
$$
and suppose $ \ J_{\beta}(K)  > 0. \ $  It is well known, see e.g. \cite{Compte}, \cite{Lerasle} that

$$
|{\bf E} f_n - f|  \le C_1 h^{\beta}\cdot J_{\beta}(K), \ \Var  \{f_n \} \ \le C_2 V_2(K)/nh.
$$
 \ The mean square error  for the considered statistics $ \ f_n \ $  allows the estimate

$$
Z_n(K) \stackrel{def}{=} E(f_n - f)^2 \le C_3 \left[ \ \frac{V_2(K)}{nh} + h^{2 \beta} J^2_{\beta}(K)  \ \right].
$$
\ The minimum  $ \ W = W(K) \ $ of the right - hand side of the last inequality relative the bandwidth $ \ h \ $ is following

 $$
 W = W(K) := \min_{h > 0} Z_n(K) \asymp C_4 \ n^{-2 \beta/(2 \beta + 1)} \ J_{\beta}(K) \cdot \left[ \ V_2(K)  \ \right]^{2 \beta}.
 $$
 \ We get to the following extremal problem relative the kernel $ \ K(\cdot) \ $ under formulated before limitations

\begin{equation} \label{Phi beta}
\Phi(K) \stackrel{def}{=} J_{\beta}(K) \cdot \left[ \ V_2(K)  \ \right]^{2 \beta} \to \min_K.
\end{equation}

\vspace{4mm}

 \ Let us apply the famous calculus of variations, see e.g. \cite{Courant Hilbert}, p.169; \cite{Gelfand Fomin},
 chapter 2, section 2.2. Namely, introduce the perturbed  kernel

$$
K_{\delta}(y) := K_0(y) + \delta \ g(y),
$$
where $ \ K_0 \ $ is the optimal kernel, $ \ \delta \ $ is "small" constant, e.g.
 $ \ -0.5 \le \delta \le 0.5, \ g = g(y), \ |y| \le  \theta = \theta_0 \ $ is suitable perturbation function. Of course,

\begin{equation} \label{perturb}
\int_{-\theta}^{\theta} g(y) \ dy = 0,
\end{equation}
"centering" condition. \par

 \ We obtain after some calculations

$$
\Phi \left(K_{\delta} \right) =  \Phi \left(K_{0} \right) +
$$

$$
C_2 \delta \int_{-\theta}^{\theta} \left\{ \ C_3 K_0(y) + C_4 |y|^{\beta} \ \right\} \ g(y) \ dy + 0(\delta^2), \ \delta \to 0.
$$

 \ Therefore, for some  finite constants $ \ C_3, C_4 \ $ and for arbitrary perturbation "centered" function $ \ g = g(y) \ $
 (\ref{perturb})

\begin{equation} \label{int o}
 \int_{-\theta}^{\theta} \left\{ \ C_3 K_0(y) + C_4 |y|^{\beta} \ \right\} \ g(y) \ dy = 0.
\end{equation}

 \ It follows  from  (\ref{int o}) taking into account the "centering" condition (\ref{perturb}) that

\begin{equation} \label{form of K}
 C_3 K_0(y) + C_4 |y|^{\beta} = C_5.
\end{equation}

 \ We conclude that the considered in this section optimization problem quite coincides with considered in the third section! \par

\vspace{3mm}

 \ Thus, the optimal kernel $ \ K_0 \ $  in this statement problem is described completely in the theorem 3.1. \par

\vspace{5mm}

\section{Estimation of the derivatives for density.}

\vspace{5mm}

  \hspace{3mm} It is interest in our opinion to find the optimal kernels for the problem of {\it derivative
 density estimations,} in the spirit, for example, \cite{Hansen}, pp. 12 - 16; where are described also some applications.\par

\vspace{3mm}

 \ Denote

$$
f^{(r)}(x) := \frac{d^r f(x)}{dx^r}, \ r = 1,2,\ldots;
$$
and we want to build the kernel estimation $ \ f_n^{(r)}(x) \ $ for the derivative $ \ f^{(r)}. \ $  \par

  \ We suppose that  for some $ \ m = 1,2,\ldots \ $ the density function $ \ f(\cdot) \ $  is $ \ r + 2m \ $  times continuous bounded
differentiable:

$$
\sup_{x \in R} | \ f^{(r + 2m)}(x) \ | < \infty.
$$

 \ As for the kernel $ \ K(\cdot); \ $ we assume in addition to the foregoing restrictions
  that it belongs to the Sobolev's space $ \ W_{2,r}(-\theta,\theta): \ $

\begin{equation} \label{Sobolev}
V_{r,2}(K) = \int_{\theta}^{\theta} \left[ \  K^{(r)}(y)  \ \right]^2 \ dy < \infty,
\end{equation}
 and as ordinary

$$
|y| \ge \theta \ \Rightarrow K(y) = 0; \ \int_{-\theta}^{\theta} K(y) \ dy = 1;
$$

$$
\forall s =1,2,\ldots, 2 m - 1 \ \Rightarrow \int_{-\theta}^{\theta} y^s \ K(y) \ dy = 0;
$$

$$
\int_{-\theta}^{\theta} y^{2m} \ K(y) \ dy = 1; \ K(-y) = K(y).
$$

\vspace{3mm}

 \ The kernel estimate for the derivative $ \ f^{(r)}(x) \ $ has a form

\begin{equation} \label{ker der}
f_n^{(r)}(x) = \frac{1}{n \ h^{1 + r}} \ \sum_{i=1}^n K^{(r)} \left( \ \frac{\xi_i - x}{h}  \ \right),
\end{equation}
where as before $ \ n \to \infty \ \Rightarrow h = h(n) \to 0, \ n h^{1 + r} \to \infty.  \ $\par

 \ It is known, see \cite{Hansen}, p.11 - 15 that the bias as $ \ n to \infty \ $ of $ \ f_n^{(r)}(x) \ $ under
 our condition has a form

$$
{\bf E} f_n^{(r)}(x) - f(x) \sim  C_1(f) \ h^{2m} \ \int_{-\theta}^{\theta} y^{2m} \ K(y) \ dy,
$$
and the variance may be evaluated as follows

$$
\Var \left[ \ f_n^{(r)}(x)  \ \right] \sim C_2(f) \ \frac{1}{n h^{1 + r}} \ \left[ \ V_{r,2}(K)  \ \right].
$$

\vspace{3mm}

 \ We get as before to the following extremal problem under our conditions

\begin{equation} \label{der opt}
\Phi(K) := \int_{-\theta}^{\theta} \left[ \ K^{(r)}(y)  \ \right]^2 \ dy \to \min.
\end{equation}

\vspace{3mm}

 \ The solution of this problem is quite alike to one in the second section. \par

\vspace{4mm}

 \ {\bf Theorem 5.1.} The optimal kernel $ \ K_r(y) \ $ for the considered in this section is unique and has a form

\begin{equation} \label{der ker optim}
K_r(y) = \frac{1}{2 \theta} - \frac{1}{2 \theta} P_{2m + 2r} \left( \ \frac{y}{\theta} \ \right), \ |y| \le \theta,
\end{equation}

$ K_r(y) = 0, \ |y| > \theta,  \ $ where

\begin{equation} \label{theta opt der}
\theta = [1 - \mu(2m + 2r)]^{-1/(2m + 2 r)}.
\end{equation}

 \vspace{5mm}

\section{Concluding remarks.}

 \vspace{5mm}

 \hspace{3mm} {\bf A.} The multivariate version of the kernel, in particular, optimal one, has a factorizable form

$$
V(x_1,x_2,\ldots,x_d) = \prod_{j=1}^d  K(x_j), \ d = 2,3,\ldots,
$$
see \cite{Bertin1}, \cite{Bertin2}, \cite{Devroe}, \cite{Epanechnikov}, \cite{Ibragimov Khasminskii 1},  \cite{Ibragimov Khasminskii 2},
\cite{Lerasle} etc. \par

\vspace{3mm}

 \ {\bf B.} At the same optimization problem for density measurement appears for the  so - called {\it recursive} Wolverton - Wagner's density
 estimation,  see \cite{Devroe}, \cite{Khardani}, \cite{Nadaraya}, \cite{Slaoui}, \cite{Wegman}, \cite{Wolverton Wagner} and so one. \par

\vspace{3mm}

 \ {\bf C.}  Offered here method may be generalized perhaps on the so called regression problem, i.e. when

$$
\eta_i = f(x_i) + \epsilon_i, \ i = 1,2,\ldots,n;
$$
see \cite{Gasser}, \cite{Hardle}, p.64, Theorem 3.1.\par

\vspace{3mm}

 \ {\bf D.} The case when the r.v. $ \ \xi_i \ $ are positive, may be reduced to the considered here by a transform

$$
\eta_i := \ln \xi_i,
$$
therefore

$$
f_{\eta}(x) = e^x \ f_{\xi} \left( e^x \right), \ x \in (-\infty, \infty),
$$
see \cite{Charpentier}. The case when $ \ \xi_i \in (a,b) \ $ may be considered quite analogously. \par

\vspace{6mm}

\vspace{0.5cm} \emph{Acknowledgement.} {\footnotesize The first
author has been partially supported by the Gruppo Nazionale per
l'Analisi Matematica, la Probabilit\`a e le loro Applicazioni
(GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM) and by
Universit\`a degli Studi di Napoli Parthenope through the project
\lq\lq sostegno alla Ricerca individuale\rq\rq (triennio 2015 - 2017)}.\par

\vspace{4mm}

 \ The second author is grateful to  Yousri Slaoui for sending its a very interest
 article  \cite{Khardani}. \par

\end{document}